Find the slope of the line that passes through the points.

(3, 5), (5, 6)
1
Expert's answer
2020-10-25T18:42:10-0400

To get the Slope (gradient) ;


"\frac{y_2-y_1}{x_2-x_1}"

="\frac{6-5}{5-3}" ="\frac{1} {2}"


Hence the Slope is "\frac {1} {2}"